Let f(x) = 4x + 4 and g(x) = 522 + 5x. After simplifying, (fog)(x) =
Answer #1

f(x)=4x+4,g(x)=5x^2+5x

f.o.g(x)=f(g(x))

We know, to find out f(g(x)) , we have to substitute the function g(x) in place of the variable 'x' in the function f(x) .

That is

(f.o.g)(x)=f(5x^2+5x)=4(5x^2+5x)+4=20x^2+20x+4

Hence, we got the simplified form